If I have 100ml of a 6% solution; how do I dilute that to a 3% solution?
Thank you,
chere
my thought is that I would take 94% of 100ml and then add that to the 6% solution?

Answer by Theo(13342)

it appears that if you add 100 ml of a 0% solution, then you will have 200 ml of a 3% solution.

6% solution in 100 ml is equal to 6 ml because 6 ml divided by 100 ml = .06 which is equivalent to 6%.

6 ml divided by 200 ml is equal to .03 which is equivalent to 3%.

the formula that can be used is:

x = the amount of 0% solution you need to add.

you get .06 * 100 + 0 * x = .03 * (100 + x)

you are adding 100 ml of 6% solution to x ml of 0% solution to get a total of (100 + x) ml of 3% solution.

the formula becomes:

.06 * 100 + 0 * x = .03 * (100 + x)

simplify to get:

.06 * 100 + 0 * x = .03 * 100 + .03 * x

simplify further to get:

6 = 3 + .03 * x

subtract 3 from both sides of the equation to get:

3 = .03 * x

divide both sides of the equation by .03 to get:

3 /.03 = x

simplify to get:

100 = x.

you need to add 100 ml of 0% solution to 100 ml of 6% solution to get 200 ml of 3% solution.